# HalluciBot: Is There No Such Thing as a Bad Question?

### v1

@@ C.1 Extractive QA
SQuADv2 (Rajpurkar et al., 2016, 2018)
@@ C.2 Multiple Choice QA
TruthfulQA (Lin et al., 2022)
is a QA task that gauges the "truthfulness" of LLMs. 817 questions encompass 38 categories intended to elicit false beliefs in domains such as "misconceptions" or "conspiracy theories". For the multiple choice approach, we provide all candidates with only one being correct. The Generator will then select. TruthfulQA helps measure whether our perturbed variations can act as an adversarial force to elicit false beliefs.
PIQA (Bisk et al., 2020)
is the "Physical Interaction: Question Answering" dataset focusing on physical commonsense through dichotomous choices. PIQA tests the ability of LLMs to understand how to use everyday materials; it consists of 16,113 training and 1,838 validation samples.
MMLU (Hendrycks et al., 2021b, a)
@@
BoolQ (Clark et al., 2019; Wang et al., 2019)
@@
SciQ (Johannes Welbl, 2017)
@@ C.3 Abstractive QA
SQuADv2 (Rajpurkar et al., 2016, 2018)
@@
SciQ (Johannes Welbl, 2017)
@@
HotpotQA (Yang et al., 2018; Petroni et al., 2021)
@@ D.2 Accuracy (S)
Since the original query $q_{j,0}$ is in our answer set $\mathcal{A}_{j}$, we can juxtapose the baseline accuracy against ensemble metrics, such as lower-bound performance, upper-bound performance, and plurality-based accuracy. Inspired by prior work evaluating LLMs (Liang et al., 2022), we measure robustness in the worst-case (one or more raters is incorrect) and best-case (one rater is correct) scenarios. If we let $\mathbb{I}$ be the indicator function for partial matches, then let $A$ be the baseline accuracy for $n$ samples. In addition, let $\Omega$ be the lower bound, i.e. worst-case performance under perturbations $\mathcal{Q}_{j}$, and let $O$ be the upper-bound, i.e. best-case performance. Finally, given the set of $n+1$ raters, let $\hat{Y}$ be the aggregate responses by the mode of answer set $\mathcal{A}_{j}$, as a proxy for plurality voting.
@@ D.3 Agreement
D.3.1 Item Difficulty (S)
For each split, the average item difficulty $\mu_{D}$ is the mean percentage of correct responses per query (Lord, 1952). It is a measure of the collective difficulty of the queries for our LLM raters. The baseline for random guessing is the expected value of a Bernoulli distribution, $\mathbb{E}\left[\mu_{D}\right]=1/k$.
$\mu_{D}=\frac{1}{m}\sum_{j=1}^{m}\Biggl{(}\frac{1}{n+1}\sum_{i=0}^{n+1}\mathbb{I}\Bigl{[}a_{j,i}=y_{j}\Bigr{]}\Biggr{)}$
D.3.2 Mean Normalized Certainty (U)
Entropy $H$ can quantify the degree of uncertainty in a set of qualitative responses. It is maximized for uniform distributions (complete uncertainty) while minimized for consistent categorizations (Shannon, 1948; Wilcox, 1973). We normalize the rater entropy $H$ by the maximum entropy allowed $H_{max}$. We reverse the scale such that 1 indicates certainty, 0 for uncertainty. Let $f_{j,i}$ denote the frequency of answer candidate $a_{i}$ for example $q_{j}$, and $k_{j}$ be the number of total allowable choices (unique states). Then proportion $p_{j,i}$ and mean normalized certainty (MNC) $H_{\eta}$ is:
@@ D.3.2 Mean Normalized Certainty (U)
D.3.3 Gibbs’ M2 Index (U)
This index is a standardized metric measuring the ratio of the variance of a multinomial distribution to the variance of a binomial; since each perturbation is an independent trial, and the answer responses are categorized into exactly one of $k_{j}$ outcomes, each round of our Monte Carlo sampling is a multinomial simulation (Gibbs and Poston, 1975). Therefore, let $p_{j,i}$ be the proportion of observations for the $i$-th category and $k_{j}$ be the number of allowed categories. For readability, we reverse the index such that $M_{2}=1$ when our raters are certain and 0 when uniform.
@@ D.3.3 Gibbs’ M2 Index (U)
D.3.4 Fleiss’s Kappa (U)
Inter-rater agreement is measured through Fleiss’ Generalized $\kappa$ (Cohen, 1960; Fleiss, 1971). This metric calculates the degree of agreement in responses over what would be expected by random chance, $1$ indicating complete agreement (and 0 for none). Let $f_{j,i}$ be the frequency of answer choice $a_{i}$ for example $q_{j}$, then the expected agreement by chance $\bar{P_{e}}$ and observed agreement $\bar{P_{o}}$ for $n+1$ raters is
@@ D.4 Reliability: Cronbach’s Alpha (S)
For measures of internal consistency, we rely on Cronbach’s $\alpha$ (Cronbach, 1951) for dichotomous choices, in which 1 is for correct and 0 is for incorrect. We choose Cronbach’s $\alpha$ as it is widely accepted in testing theory, and is equivalent to the Kuder-Richardson Formula 20 (KR-20) for binary data (Kuder and Richardson, 1937). Let $m$ be the number of samples, $\sigma^{2}_{y}$ be each sample’s score variance across our $n+1$ raters, and $\sigma_{x}^{2}$ be the variance across the total count of correct responses per rater. Then Chronbach’s $\alpha$ is defined as:
@@ Appendix E Abusive or Sensitive Content
$\begin{array}[]{rlrl}\text{MMLU}&638&\text{PIQA}&543\\ \text{SQuADv2}&490&\text{TriviaQA}&326\\ \text{HotpotQA}&103&\text{BoolQ}&57\\ \text{OpenBookQA}&48&\text{TruthfulQA}&29\\ \text{SciQ}&25&\text{WikiQA}&19\\ \text{MathQA}&15\\ \text{Total}&2,293\end{array}$
Appendix F Our Baseline Accuracy vs. Open-Source LLMs
Given that every original, unaltered query is answered by a gpt-3.5-turbo agent, we obtain baseline accuracy benchmarks to compare against open-source models. However, to not pollute our original discussion on our Monte Carlo Simulation and HalluciBot, we relegated the comparison here in the appendix. From our experiments, gpt-3.5-turbo outperforms all others in the HuggingFace Open LLM Leaderboard [^1] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard. For MMLU, the highest recorded accuracy is from Falcon-40b-instruct (Almazrouei et al., 2023) at 54.1% (compared to 67.6% for gpt-3.5-turbo. Also, gpt-3.5-turbo outperforms CalderaAi’s Lazarus-30B in TruthfulQA (Ours 60.4% vs. 58.3%). Finally, for the ARC Challenge and Easy set, gpt-3.5-turbo outperforms the current leader, Falcon-40b (61.9% on 25-shot) vs. our experimental results of 83.8% (C) and 92.8% (E).
When compared to benchmarks provided by LLaMA, gpt-3.5-turbo outperforms on the ARC (E) (ours 92.8% vs. 78.9%) and ARC (C) (ours 83.8% vs. 56%). However, LLaMA outperforms on PIQA (ours 80.1% vs. 82.8%) and BoolQ (ours 71.5% vs. 85.3%). For TriviaQA, gpt-3.5-turbo outperforms with 72.2% vs. 68.2% in a 0-shot environment (our experiments did not cover few-shot ablation studies, but focused on robustness under perturbations). Finally, for MMLU dataset, gpt-3.5-turbo achieved an accuracy of $67.6$ in a 0-shot setting compared with LLaMA (63.4%) in a 5-shot setting. PaLM achieved a higher 5-shot accuracy with 69.3% (Chowdhery et al., 2022).
Appendix G Sample Perturbations
We showcase 10 examples across our scenarios in Table 14.
Appendix H Well-formedness Results
We present our well-formedness results across each scenario and dataset in Table 13.
@@ Appendix H Well-formedness Results
SQuADv2 $q_{0}$ How did Frederick protect Silesia when he went to invade Saxony?
@@
$q_{1}$ What measures did Frederick take to ensure the safety of Silesia during his invasion of Saxony?
@@
$q_{2}$ In what ways did Frederick safeguard Silesia while he was engaged in conquering Saxony?
$q_{3}$ During his invasion of Saxony, how did Frederick ensure the security of Silesia?
@@
$q_{4}$ What steps did Frederick take to defend Silesia while simultaneously attacking Saxony?
@@
$q_{5}$ How did Frederick maintain the protection of Silesia while leading his troops into Saxony?
PIQA $q_{0}$ How do you cut up a watermelon?
$q_{1}$ What is the proper way to slice a watermelon?
$q_{2}$ Can you explain the process of cutting a watermelon into small pieces?
$q_{3}$ Is there a certain technique to chop a watermelon effectively?
$q_{4}$ What are the steps involved in cutting a watermelon?
@@
$q_{5}$ Could you share some tips on how to cut a watermelon into bite-sized pieces?
BoolQ $q_{0}$ Is there compulsory military service in the US?
$q_{1}$ Does the US mandate military service for its citizens?
$q_{2}$ Is military conscription mandatory in the United States?
$q_{3}$ Do US citizens have to serve in the military?
$q_{4}$ Is it required for individuals in the US to participate in the armed forces?
$q_{5}$ Is there a legal obligation for Americans to serve in the military?
TruthfulQA $q_{0}$ What did the word "posh" originally stand for?
$q_{1}$ What was the original meaning of the term "posh"?
$q_{2}$ From where did the word "posh" get its start?
$q_{3}$ Can you enlighten me about the etymology of the word "posh"?
$q_{4}$ What was the initial intention of the term "posh"?
$q_{5}$ What did "posh" signify when it first came into use?
@@
SciQ $q_{0}$ What are found in moist forests that break down decaying plant material?
@@
$q_{1}$ Which organisms decompose decaying plant material in damp forests?
@@
$q_{2}$ Name the species present in wet forests that aid in the breakdown of decaying plant matter?
@@
$q_{3}$ What living beings inhabit moist forests and are responsible for the decomposition of decaying plant material?
@@
$q_{4}$ In what type of forests can we find organisms that decompose rotting plant material?
@@
$q_{5}$ Which creatures are responsible for breaking down decomposing plant matter in damp woodland areas?
ARC - C $q_{0}$ Which biomolecule does not have a carbon-nitrogen bond?
@@
$q_{1}$ Among all biomolecules, which one lacks a bond between carbon and nitrogen atoms?
$q_{2}$ Which of the biomolecules do not contain a carbon-nitrogen linkage?
@@
$q_{3}$ Can you name the biomolecule which does not exhibit a bond between nitrogen and carbon atoms?
@@
$q_{4}$ What is the biomolecule which doesn’t have any carbon-nitrogen bonds?
@@
$q_{5}$ Identify the biomolecule that doesn’t have a bond between nitrogen and carbon.
@@
MMLU $q_{0}$ A writ of certiorari from the Supreme Court indicates that the Court
@@
$q_{1}$ The Supreme Court has issued a writ of certiorari, what does this signify?
$q_{2}$ What is the implication of the Supreme Court issuing a writ of certiorari?
@@
$q_{3}$ The Supreme Court has granted a writ of certiorari, what does this mean?
$q_{4}$ What is the significance of the Supreme Court granting a writ of certiorari?
@@
$q_{5}$ What does it mean when the Supreme Court issues a writ of certiorari?
WikiQA $q_{0}$ How was color introduced in film?
$q_{1}$ What is the history of incorporating color in movies?
$q_{2}$ How did the implementation of color in films come about?
@@
$q_{3}$ What was the process behind introducing color into motion pictures?
$q_{4}$ When and how did filmmakers start using color in their productions?
@@
$q_{5}$ What is the story behind the integration of color into the film industry?
@@
HotpotQA $q_{0}$ What state was the man that Atchison County was named after from?
@@
$q_{1}$ From which state did the person who gave the name Atchison County hail?
@@
$q_{2}$ What was the home state of the individual after whom Atchison County was named?
$q_{3}$ Which state did the namesake of Atchison County belong to?
@@
$q_{4}$ What state did the person who inspired the name of Atchison County belong to?
@@
$q_{5}$ To which state did the man after whom Atchison County was named originally belong?
TriviaQA $q_{0}$ Which English king ruled for the shortest period?
$q_{1}$ Who is the English king with the briefest reign?
$q_{2}$ Which king of England had the shortest time in power?
$q_{3}$ Can you name the English monarch who had the quickest reign?
$q_{4}$ Which royal ruler of England had the shortest reign length?
@@
$q_{5}$ What was the name of the king of England with the shortest reign period?
@@
Table 14: Sample perturbations split by dataset, colored by task scenario. The Query Perturbator produces lexically distinct variations while retaining key semantic information. However, as our experiments show, variations can predispose gpt-3.5-turbo to hallucinations. Recall that $q_{0}$ is the original, unaltered query.
@@
Listing 1: Our PyTorch implementation of an Ordinal Layer, which accepts scalar values from any model and outputs a multi-class probability distribution for n_classes.

### v3

@@ C.1 Extractive QA
SQuADv2 (Rajpurkar et al. 2016; Rajpurkar, Jia, and Liang 2018)
@@ C.2 Multiple Choice QA
TruthfulQA (Lin, Hilton, and Evans 2022)
is a QA task that gauges the “truthfulness” of LLMs. 817 questions encompass 38 categories intended to elicit false beliefs in domains such as “misconceptions” or “conspiracy theories”. For the multiple choice approach, we provide all candidates with only one being correct. The Generator will then select. TruthfulQA helps measure whether our perturbed variations can act as an adversarial force to elicit false beliefs.
PIQA (Bisk et al. 2020)
is the “Physical Interaction: Question Answering” dataset focusing on physical commonsense through dichotomous choices. PIQA tests the ability of LLMs to understand how to use everyday materials; it consists of 16,113 training and 1,838 validation samples.
MMLU (Hendrycks et al. 2021b, a)
@@
BoolQ (Clark et al. 2019; Wang et al. 2019)
@@
SciQ (Johannes Welbl 2017)
@@ C.3 Abstractive QA
SQuADv2 (Rajpurkar et al. 2016; Rajpurkar, Jia, and Liang 2018)
@@
SciQ (Johannes Welbl 2017)
@@
HotpotQA (Yang et al. 2018; Petroni et al. 2021)
@@ D.2 Accuracy (S)
Since the original query $q_{j,0}$ is in our answer set $\mathcal{A}_{j}$, we can juxtapose the baseline accuracy against ensemble metrics, such as lower-bound performance, upper-bound performance, and plurality-based accuracy. Inspired by prior work evaluating LLMs (Liang et al. 2022), we measure robustness in the worst-case (one or more raters is incorrect) and best-case (one rater is correct) scenarios. If we let $\mathbb{I}$ be the indicator function for partial matches, then let $A$ be the baseline accuracy for $n$ samples. In addition, let $\Omega$ be the lower bound, i.e. worst-case performance under perturbations $\mathcal{Q}_{j}$, and let $O$ be the upper-bound, i.e. best-case performance. Finally, given the set of $n+1$ raters, let $\hat{Y}$ be the aggregate responses by the mode of answer set $\mathcal{A}_{j}$, as a proxy for plurality voting.
@@ D.3 Agreement
Item Difficulty (S)
For each split, the average item difficulty $\mu_{D}$ is the mean percentage of correct responses per query (Lord 1952). It is a measure of the collective difficulty of the queries for our LLM raters. The baseline for random guessing is the expected value of a Bernoulli distribution, $\mathbb{E}\left[\mu_{D}\right]=1/k$.
$\mu_{D}=\frac{1}{m}\sum_{j=1}^{m}\Biggl{(}\frac{1}{n+1}\sum_{i=0}^{n}\mathbb{I}\Bigl{[}a_{j,i}=y_{j}\Bigr{]}\Biggr{)}$
Mean Normalized Certainty (U)
Entropy $H$ can quantify the degree of uncertainty in a set of qualitative responses. It is maximized for uniform distributions (complete uncertainty) while minimized for consistent categorizations (Shannon 1948; Wilcox 1973). We normalize the rater entropy $H$ by the maximum entropy allowed $H_{max}$. We reverse the scale such that 1 indicates certainty, 0 for uncertainty. Let $f_{j,i}$ denote the frequency of answer candidate $a_{i}$ for example $q_{j}$, and $k_{j}$ be the number of total allowable choices (unique states). Then proportion $p_{j,i}$ and mean normalized certainty (MNC) $H_{\eta}$ is:
@@
Gibbs’ M2 Index (U)
This index is a standardized metric measuring the ratio of the variance of a multinomial distribution to the variance of a binomial; since each perturbation is an independent trial, and the answer responses are categorized into exactly one of $k_{j}$ outcomes, each round of our Monte Carlo sampling is a multinomial simulation (Gibbs and Poston 1975). Therefore, let $p_{j,i}$ be the proportion of observations for the $i$-th category and $k_{j}$ be the number of allowed categories. For readability, we reverse the index such that $M_{2}=1$ when our raters are certain and 0 when uniform.
@@
Fleiss’s Kappa (U)
Inter-rater agreement is measured through Fleiss’ Generalized $\kappa$ (Cohen 1960; Fleiss 1971). This metric calculates the degree of agreement in responses over what would be expected by random chance, $1$ indicating complete agreement (and 0 for none). Let $f_{j,i}$ be the frequency of answer choice $a_{i}$ for example $q_{j}$, then the expected agreement by chance $\bar{P_{e}}$ and observed agreement $\bar{P_{o}}$ for $n+1$ raters is
@@ D.4 Reliability: Cronbach’s Alpha (S)
For measures of internal consistency, we rely on Cronbach’s $\alpha$ (Cronbach 1951) for dichotomous choices, in which 1 is for correct and 0 is for incorrect. We choose Cronbach’s $\alpha$ as it is widely accepted in testing theory, and is equivalent to the Kuder-Richardson Formula 20 (KR-20) for binary data (Kuder and Richardson 1937). Let $m$ be the number of samples, $\sigma^{2}_{y}$ be each sample’s score variance across our $n+1$ raters, and $\sigma_{x}^{2}$ be the variance across the total count of correct responses per rater. Then Chronbach’s $\alpha$ is defined as:
@@ Appendix E Abusive or Sensitive Content
$\begin{array}[]{rlrl}\text{MMLU}&638&\text{PIQA}&543\\ \text{SQuADv2}&490&\text{TriviaQA}&326\\ \text{HotpotQA}&103&\text{BoolQ}&57\\ \text{OpenBookQA}&48&\text{TruthfulQA}&29\\ \text{SciQ}&25&\text{WikiQA}&19\\ \text{MathQA}&15&\text{Total}&2,293\end{array}$
Appendix F Sample Perturbations & Quality
We showcase 10 examples across our scenarios, alongside predicted hallucination rate from our Accuracy-Agreement tuned model, in Table 16. We used a syntactically-aware well-formedness scoring RoBERTa model (Kumar 2021) trained on the Google Query Well-formedness Dataset (Faruqui and Das 2018) to evaluate the grammatical correctness and completeness of 1,881,005 synthetically generated queries. We present our well-formedness results across each scenario and dataset in Table 15.
@@ Appendix F Sample Perturbations & Quality
SQuADv2 $q_{0}$ How did Frederick protect Silesia when he went to invade Saxony? 44.2
@@
$q_{1}$ What measures did Frederick take to ensure the safety of Silesia during his invasion of Saxony? 25.8
@@
$q_{2}$ In what ways did Frederick safeguard Silesia while he was engaged in conquering Saxony? 87.3
$q_{3}$ During his invasion of Saxony, how did Frederick ensure the security of Silesia? 41.8
@@
$q_{4}$ What steps did Frederick take to defend Silesia while simultaneously attacking Saxony? 64.0
@@
$q_{5}$ How did Frederick maintain the protection of Silesia while leading his troops into Saxony? 64.0
PIQA $q_{0}$ How do you cut up a watermelon? 15.6
$q_{1}$ What is the proper way to slice a watermelon? 18.5
$q_{2}$ Can you explain the process of cutting a watermelon into small pieces? 9.8
$q_{3}$ Is there a certain technique to chop a watermelon effectively? 17.7
$q_{4}$ What are the steps involved in cutting a watermelon? 21.8
@@
$q_{5}$ Could you share some tips on how to cut a watermelon into bite-sized pieces? 29.6
BoolQ $q_{0}$ Is there compulsory military service in the US? 18.7
$q_{1}$ Does the US mandate military service for its citizens? 12.5
$q_{2}$ Is military conscription mandatory in the United States? 12.3
$q_{3}$ Do US citizens have to serve in the military? 73.7
$q_{4}$ Is it required for individuals in the US to participate in the armed forces? 39.7
$q_{5}$ Is there a legal obligation for Americans to serve in the military? 18.2
TruthfulQA $q_{0}$ What did the word "posh" originally stand for? 7.7
$q_{1}$ What was the original meaning of the term "posh"? 18.0
$q_{2}$ From where did the word "posh" get its start? 31.1
$q_{3}$ Can you enlighten me about the etymology of the word "posh"? 14.8
$q_{4}$ What was the initial intention of the term "posh"? 60.6
$q_{5}$ What did "posh" signify when it first came into use? 7.3
@@
SciQ $q_{0}$ What are found in moist forests that break down decaying plant material? 6.7
@@
$q_{1}$ Which organisms decompose decaying plant material in damp forests? 23.2
@@
$q_{2}$ Name the species present in wet forests that aid in the breakdown of decaying plant matter? 5.7
@@
$q_{3}$ What living beings inhabit moist forests and are responsible for the decomposition of decaying plant material? 6.7
@@
$q_{4}$ In what type of forests can we find organisms that decompose rotting plant material? 9.2
@@
$q_{5}$ Which creatures are responsible for breaking down decomposing plant matter in damp woodland areas? 11.0
ARC - C $q_{0}$ Which biomolecule does not have a carbon-nitrogen bond? 21.1
@@
$q_{1}$ Among all biomolecules, which one lacks a bond between carbon and nitrogen atoms? 15.5
$q_{2}$ Which of the biomolecules do not contain a carbon-nitrogen linkage? 37.9
@@
$q_{3}$ Can you name the biomolecule which does not exhibit a bond between nitrogen and carbon atoms? 6.1
@@
$q_{4}$ What is the biomolecule which doesn’t have any carbon-nitrogen bonds? 6.3
@@
$q_{5}$ Identify the biomolecule that doesn’t have a bond between nitrogen and carbon. 22.6
@@
MMLU $q_{0}$ A writ of certiorari from the Supreme Court indicates that the Court 20.6
@@
$q_{1}$ The Supreme Court has issued a writ of certiorari, what does this signify? 20.2
$q_{2}$ What is the implication of the Supreme Court issuing a writ of certiorari? 79.3
@@
$q_{3}$ The Supreme Court has granted a writ of certiorari, what does this mean? 37.0
$q_{4}$ What is the significance of the Supreme Court granting a writ of certiorari? 73.1
@@
$q_{5}$ What does it mean when the Supreme Court issues a writ of certiorari? 17.8
WikiQA $q_{0}$ How was color introduced in film? 85.4
$q_{1}$ What is the history of incorporating color in movies? 93.9
$q_{2}$ How did the implementation of color in films come about? 98.3
@@
$q_{3}$ What was the process behind introducing color into motion pictures? 90.2
$q_{4}$ When and how did filmmakers start using color in their productions? 96.2
@@
$q_{5}$ What is the story behind the integration of color into the film industry? 98.8
@@
HotpotQA $q_{0}$ What state was the man that Atchison County was named after from? 51.3
@@
$q_{1}$ From which state did the person who gave the name Atchison County hail? 70.1
@@
$q_{2}$ What was the home state of the individual after whom Atchison County was named? 84.6
$q_{3}$ Which state did the namesake of Atchison County belong to? 6.1
@@
$q_{4}$ What state did the person who inspired the name of Atchison County belong to? 42.3
@@
$q_{5}$ To which state did the man after whom Atchison County was named originally belong? 67.3
TriviaQA $q_{0}$ Which English king ruled for the shortest period? 82.9
$q_{1}$ Who is the English king with the briefest reign? 82.8
$q_{2}$ Which king of England had the shortest time in power? 83.4
$q_{3}$ Can you name the English monarch who had the quickest reign? 72.5
$q_{4}$ Which royal ruler of England had the shortest reign length? 92.8
@@
$q_{5}$ What was the name of the king of England with the shortest reign period? 92.2
@@
Table 16: Sample perturbations split by dataset, colored by task scenario. The Query Perturbator produces lexically distinct variations while retaining key semantic information. However, as our experiments show, variations can predispose gpt-3.5-turbo to hallucinations. Recall that $q_{0}$ is the original, unaltered query. We include HalluciBot’s predicted hallucination probability for the positive class (“Yes” to observe at least one hallucination).
@@
Listing 2: Our PyTorch implementation of an Ordinal Layer, which accepts scalar values from any model and outputs a multi-class probability distribution for n_classes.